# Covariance test and universal bootstrap by operator norm

### v1

@@ 3.4 Sharp uniform simultaneous confidence intervals
$\bm{c}_{1}^{T}\bm{\hat{\Sigma}}\bm{c}_{2}+\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)\ \|\bm{c}_{1}\|\cdot\|\bm{c}_{2}\|\bigg{]}\bigg{|}\bm{Y}^{1},\cdots,\bm{Y}^{B}\bigg{)}-(1-\alpha)\bigg{|}$ $\lesssim\rho_{n}(\bm{\Sigma}_{0})+d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))+B^{-\frac{1}{2}}.$
Theorem 3.6 establishes sharp uniform simultaneous confidence intervals for both $\langle\bm{A},\bm{\Sigma}\rangle$ and $\bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}$. Specifically, intervals (40) and (41) are sharp, as their probability converges uniformly to $1-\alpha$, rather than no less than $1-\alpha$. When $\bm{c}_{1}$ and $\bm{c}_{2}$ are chosen as the canonical basis vectors $\bm{e}_{1},\cdots,\bm{e}_{p}$ in $\mathbb{R}^{p}$, we obtain a uniform entry-wise confidence interval for all $\sigma_{ij}$. Notably, Corollary 3.6 extends this result by providing confidence intervals for all linear combinations of entries of $\bm{\Sigma}$. Constructing confidence intervals for all such combinations is considerably more challenging than for individual entries. For a mean vector in the simpler vector case, achieving confidence intervals for each entry requires only $\ln p=o(n^{a})$ for some $a>0$, see Chernozhukov, Chetverikov and Kato (2013, 2017); Chernozhukov, Chetverikov and Koike (2023). In contrast, confidence intervals for all linear combinations of the mean vector require $p=o(n)$, see Zhai (2018); Xu, Zhang and Wu (2019); Fang and Koike (2024). Remarkably, we are able to construct sharp simultaneous confidence intervals for all $\langle\bm{A},\bm{\Sigma}\rangle$ even as $p/n$ converges or diverges to infinity.
4 Power analysis and the generalized universal bootstrap approximation
4.1 Power analysis for operator norm-based statistics
As noted in Section 1, several popular statistics utilize the operator norm in covariance testing, such as Roy’s largest root statistic $T^{\text{Roy}}$ in (3). Our Theorem 3.5 also applies to Roy’s largest root with universal bootstrap, prompting the question of which test to use. In this subsection, we perform a power analysis across a family of statistics, with $T$ and $T^{\text{Roy}}$ examined as specific cases.
We conduct a power analysis under the generalized spike model setting, which allows variability in the bulk eigenvalues and does not assume block independence between the spiked and bulk parts, as in Bai and Yao (2012) and further developed by Jiang and Bai (2021). We assume the following generalized spike structure
where $\bm{D},\bm{V}_{1},\bm{R}_{1}\in\mathbb{R}^{k\times k}$, $\bm{V}_{2},\bm{R}_{2}\in\mathbb{R}^{(p-k)\times(p-k)}$ are diagonal matrices, $k$ is a fixed number, and $\bm{U}$ is orthogonal matrix. Thus
$\bm{\Sigma}_{1}^{-\frac{1}{2}}\bm{\Sigma}\bm{\Sigma}_{1}^{-\frac{1}{2}}=\bm{U}\left[\begin{array}[]{cc}\bm{D}^{\prime}&\bm{0}\\ \bm{0}&\bm{V}^{\prime}_{2}\end{array}\right]\bm{U}^{T},\ \bm{\Sigma}_{1}^{-\frac{1}{2}}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-\frac{1}{2}}=\bm{U}\left[\begin{array}[]{cc}\bm{V}^{\prime}_{1}&\bm{0}\\ \bm{0}&\bm{V}^{\prime}_{2}\end{array}\right]\bm{U}^{T},$ (53)
where $\bm{D}^{\prime}=\bm{D}\bm{R}^{-1}_{1}$, $\bm{V}_{1}^{\prime}=\bm{V}_{1}\bm{R}^{-1}_{1}$, $\bm{V}_{2}^{\prime}=\bm{V}_{2}\bm{R}^{-1}_{2}$. We assume that $\|\bm{D}^{\prime}\|_{\text{op}}$, $\|\bm{V}^{\prime}_{1}\|_{\text{op}}$, $\|\bm{V}^{\prime}_{2}\|_{\text{op}}$ are bounded, without requiring the eigenvalues of $\bm{V}_{2}$ to be the same, unlike the classical spike model. We consider the family of statistics
$T(\bm{\Sigma}_{1})=\left\|\bm{\Sigma}_{1}^{-\frac{1}{2}}\left(\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right)\bm{\Sigma}_{1}^{-\frac{1}{2}}\right\|_{\text{op}},$ (54)
and reject $H_{0}$ if $T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2}}(\alpha)$. Notably, $T=T(\bm{I}_{p})$, $T^{\text{Roy}}=T(\bm{\Sigma}_{0})$ are both included in this family. Define $E_{+}(\bm{\Sigma}_{1})$, $E_{-}(\bm{\Sigma}_{1})$ and $\tilde{m}(z;\bm{\Sigma}_{1})$ as in Section 3 with $(\bm{\Sigma},\bm{R})=(\bm{V}_{2}^{\prime},-\bm{V}_{2}^{\prime})$ in $\bm{M}_{n}$ in (14). Intuitively, $E_{+}(\bm{\Sigma}_{1})$, $E_{-}(\bm{\Sigma}_{1})$ and $\tilde{m}(z;\bm{\Sigma}_{1})$ are limits of the upper and lower bound and the Stieltjes transform of bulk eigenvalues of $\phi^{-1/2}\bm{\Sigma}_{1}^{-1/2}\left(\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right)\bm{\Sigma}_{1}^{-1/2}$. For simplicity, we restrict our analysis to $\alpha=1$ and $|E_{+}(\bm{\Sigma}_{1})|>|E_{-}(\bm{\Sigma}_{1})|$.
Theorem 4.1.
Under Assumptions 1 and (48), define $\bm{D}^{\prime}=\text{diag}\left((d_{i}^{\prime})_{i=1}^{k}\right)$, $\bm{V}_{1}^{\prime}=\text{diag}\left((v_{i}^{\prime})_{i=1}^{k}\right)$ and the threshold
$\kappa=-\frac{\phi^{\frac{1}{2}}}{\tilde{m}(\phi^{-\frac{1}{2}}E_{+};\bm{\Sigma}_{1})}-\frac{\phi^{\frac{1}{2}}v_{i}^{\prime}}{E_{+}\tilde{m}(\phi^{-\frac{1}{2}}E_{+};\bm{\Sigma}_{1})}.$
(i) If for some $1\leq i\leq k$ we have $d_{i}^{\prime}>\kappa$, then
$\mathbb{P}\left(T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-\frac{1}{2}}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-\frac{1}{2}}}(\alpha)\right)\to 1,\ \text{as}\ n\to\infty,\ B\to\infty,$
i.e. the power goes to $1$.
(ii) If for all $1\leq i\leq k$, we have $d_{i}^{\prime}<\kappa$ and $T(\bm{\Sigma}_{1})=\lambda_{1}\left(\bm{\Sigma}_{1}^{-1/2}\left(\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right)\bm{\Sigma}_{1}^{-1/2}\right)$, then
$\mathbb{P}\left(T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2}}(\alpha)\right)\to\alpha,\ \text{as}\ n\to\infty,\ B\to\infty,$
i.e. there is no power under this setting.
Theorem 4.1 provides the power analysis of the test based on the statistic (54). Specifically, it identifies the phase transition point
$-\frac{\phi^{\frac{1}{2}}}{\tilde{m}(\phi^{-\frac{1}{2}}E_{+};\bm{\Sigma}_{1})}-\frac{\phi^{\frac{1}{2}}v_{i}^{\prime}}{E_{+}\tilde{m}(\phi^{-\frac{1}{2}}E_{+};\bm{\Sigma}_{1})},$
which closely resembles the well-known Baik-Ben Arous-Péché (BBP) phase transition for the largest eigenvalues of the sample covariance matrix, see Baik, Ben Arous and Péché (2005); Baik and Silverstein (2006); Paul (2007). When the spike eigenvalues exceed this phase transition point, they lie outside the bulk eigenvalue support, yielding full test power. Conversely, if the spike eigenvalues $d^{\prime}_{i}$ fall below this point, distinguishing them from the null case using $T(\bm{\Sigma}_{1})$ becomes infeasible, as the spike and bulk eigenvalues are indistinguishably close. It is noted that Theorem 4.1 focuses solely on the largest eigenvalue spike, as analyzing the smallest eigenvalue spike is more complex. Since this power analysis serves as a preliminary illustration, we do not extend our investigation in that direction.
Corollary 3.
Under the same assumption in theorem 4.1, define $\bm{D}=\text{diag}\left((d_{i})_{i=1}^{k}\right)$, $\bm{V}_{1}=\text{diag}\left((v_{i})_{i=1}^{k}\right)$.
(i) If for some $1\leq i\leq k$, we have $d_{i}-v_{i}>-\frac{\phi^{1/2}}{\tilde{m}(\phi^{-1/2}E_{+})}-v_{i}\left(1+\frac{\phi^{1/2}}{E_{+}\tilde{m}(\phi^{-1/2}E_{+})}\right)$, then
$\mathbb{P}\left(T\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)\right)\to 1,\ \text{as}\ n\to\infty,\ B\to\infty,$
i.e. the power goes to $1$. Here $\tilde{m}(z)=\tilde{m}(z;\bm{\Sigma})$ satisfies $E_{+}\tilde{m}(E_{+})<-1$.
(ii) If for some $1\leq i\leq k$, we have $d_{i}-v_{i}>\phi^{1/2}v_{i}$, then
$\mathbb{P}\left(T^{\text{Roy}}\geq\hat{q}^{\text{ub},B}_{\bm{I}_{p}}(\alpha)\right)\to 1,\ \text{as}\ n\to\infty,\ B\to\infty,$
i.e. the power goes to $1$.
We compare the phase transition points (i) and (ii) for $T$ and $T^{\text{Roy}}$, respectively. These points reveal distinct behaviors: as $\phi^{-1/2}E_{+}\tilde{m}(\phi^{-1/2}E_{+})<-1$, the phase transition point of (i) decreases with $v_{i}$, whereas the phase transition point of (ii) increases with $v_{i}$. This indicates that when spikes affect the larger eigenvalues, i.e. for large $v_{i}$, then the statistics $T$ achieves full power more effectively than the normalized $T^{\text{Roy}}$ due to a lower phase transition point. Conversely, the normalized statistic $T^{\text{Roy}}$ performs better for smaller $v_{i}$. Thus, Corollary 3 quantifies this intuitive observation.
4.2 The generalized universal bootstrap
Given that neither $T$ and $T^{\text{Roy}}$ dominates across all settings, a natural approach is to combine these two statistics to enhance testing power, see $T^{\text{Com}}$ in (9) as an example. In Section 5, simulations demonstrate that $T^{\text{Com}}$ can outperform both $T$ and $T^{\text{Roy}}$ in certain scenarios. However, conducting tests with $T^{\text{Com}}$ requires calculating an appropriate threshold. While Theorem 3.5 provides thresholds for $T$ and $T^{\text{Roy}}$ individually, it cannot be directly applied to $T^{\text{Com}}$ due to the complex dependence between $T$ and $T^{\text{Roy}}$. To address this, we develop a generalized universality theorem that applies to a broad class of statistics involving extreme singular values.
We consider the statistic $T^{\text{ExS}}$ as a general function of extreme singular values
@@ 4.2 The generalized universal bootstrap
$T^{\text{ExS}}=f\bigg{(}\bigg{\{}\left(\sigma_{k}\left(\bm{\Sigma}_{2,1}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,1})\bm{\Sigma}_{2,1}^{-\frac{1}{2}}\right)\right)_{k=1}^{k_{1}},\cdots,\left(\sigma_{k}\left(\bm{\Sigma}_{2,M}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,M})\bm{\Sigma}_{2,M}^{-\frac{1}{2}}\right)\right)_{k=1}^{k_{M}}\bigg{\}}\bigg{)},$ (55)
where $f$ is a measurable function, $(\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m})$ are matrices, and $k_{m}$ denotes integers for $m=1,\cdots,M$. The universal bootstrapped version, $T^{\text{ExS,ub}}$, is defined by replacing $\bm{\hat{\Sigma}}$ with $\bm{\hat{\Sigma}}^{\text{ub}}$ in (55). We then define the empirical universal bootstrap threshold $\hat{q}^{\text{ExS,ub,B}}_{\bm{\Sigma},\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M}}(\alpha)$ as the upper $\alpha$-th quantile of $T^{\text{ExS,ub,1}},\cdots,T^{\text{ExS,ub,B}}$ of the i.i.d. samples $T^{\text{ExS,ub}}$. This setup enables the following generalized universal bootstrap consistency result.
Theorem 4.2 (Generalized universal bootstrap consistency).
Under Assumption 1, given $M$ admissible pairs $(\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}\bm{\Sigma}_{2,m}^{-1/2},\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}_{1,m}\bm{\Sigma}_{2,m}^{-1/2})$ and fixed integer numbers $k_{m}$ for $m=1,\cdots,M$, we have the uniform Gaussian approximation bound
$\rho_{n}(\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M})\leq Cn^{-\delta}\to 0\quad\text{as}\quad n\to\infty,$ (56)
for some constants $C$, $\delta>0$, where $\rho_{n}(\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M})$ is defined in (11).
Moreover, we have
Theorem 4.2 extends Theorem 3.5 by demonstrating that universal bootstrap consistency applies not only to $T$ and $T^{\text{Roy}}$, but also to combinations like $T^{\text{Com}}$, the sum of the largest $k$-th singular values of $\bm{\hat{\Sigma}}$ proposed by Ke (2016), and to all statistics based on extreme singular values with arbitrary dependencies in the form of (55). This result reinforces our universal bootstrap principle: to approximate the distribution of statistics involving extreme eigenvalues, it suffices to substitute all random variables with their Gaussian counterparts, showcasing the practical advantage of universality.
4.3 Comparison with other norms
In this subsection, we compare statistics based on the operator norm with those based on the two widely used norms: the Frobenius norm and the supremum norm. Define the statistics
$T^{\text{F}}=\left\|\bm{\Sigma}^{-\frac{1}{2}}\bm{\hat{\Sigma}}\bm{\Sigma}^{-\frac{1}{2}}-\bm{I}_{p}\right\|_{\text{\text{F}}},\ T^{\text{sup}}=\left\|\bm{\Sigma}^{-\frac{1}{2}}\bm{\hat{\Sigma}}\bm{\Sigma}^{-\frac{1}{2}}-\bm{I}_{p}\right\|_{\text{\text{sup}}}.$ (58)
While the universality Theorem 3.3 holds for operator norm-based statistics $T$ and $T^{\text{Roy}}$, we demonstrate that it does not extend to $T^{\text{F}}$ or $T^{\text{sup}}$. Theorem 3.3 specifies that the asymptotic distribution of $T$ and $T^{\text{Roy}}$ depends solely on the first two moments of $\bm{X}$, allowing for the construction of a universal bootstrap procedure. In contrast, the asymptotic distributions of $T^{\text{F}}$ and $T^{\text{sup}}$ rely on the first four moments of $\bm{X}$. We present the following modified assumptions to establish the asymptotic distribution of $T^{\text{F}}$.
Assumption 1.
Suppose that $\bm{Z}=\left\{Z_{ij},i=1,\cdots,n;j=1,\cdots,p\right\}$ are i.i.d. random variables with $\mathbb{E}[Z_{ij}]=0$, $\mathbb{E}[Z_{ij}^{2}]=1$, $\mathbb{E}[Z_{ij}^{4}]=\nu_{4}$. Assume that $\mathbb{E}[|Z_{ij}|^{6+\epsilon}]<\infty$ for some $\epsilon>0$.
Proposition 1.
Under Assumption 1, if $\alpha=1$ or $\alpha\geq 2$, we have
$\frac{n}{p}(T^{\text{F}})^{2}-\frac{1}{p}\left(\text{tr}(\bm{\hat{\Sigma}}\bm{\Sigma}^{-1})\right)^{2}-(\nu_{4}-2)\Rightarrow\mathcal{N}(0,4).$
To facilitate the Gaussian approximation of $T^{\text{sup}}$, we introduce the following sub-Gaussian assumption. This assumption is employed here for simplicity and can be extended to a more general, though tedious bound if needed.
Assumption 1′′.
Suppose that for $i=1,\cdots,n;\ j_{1},j_{2}=1,\cdots,p$,
$\mathbb{E}\left[\text{exp}(|Z_{ij_{1}}Z_{ij_{2}}|/B_{n})\right]\leq 2,$
for some positive sequence $B_{n}$.
Assumption 2′′.
We have for $i=1,\cdots,n;\ j_{1},j_{2}=1,\cdots,p$,
$b_{1}^{2}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}(Z_{ij_{1}}Z_{ij_{2}}-\sigma_{j_{1}j_{2}})^{2}\big{]},\ \frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}(Z_{ij_{1}}Z_{ij_{2}}-\sigma_{j_{1}j_{2}})^{4}\big{]}\leq B_{n}^{2}b_{2}^{2}$
for some positive numbers $b_{1}$, $b_{2}$ and sequence $B_{n}$ in Assumption 1′′.
Proposition 2.
If Assumptions 1′′, 2′′ hold, we have
$\sup_{t\geq 0}\bigg{|}\mathbb{P}\bigg{(}T^{\text{sup}}\leq t\bigg{)}-\mathbb{P}\bigg{(}\|\bm{G}\|_{\text{sup}}\leq t\bigg{)}\bigg{|}\lesssim\left(\frac{B_{n}^{2}\text{ln}^{2}(pn)}{n}\right)^{\frac{1}{4}},$
where $\bm{G}\in\mathbb{R}^{p\times p}$ is a Gaussian random matrix with zero mean and covariance $\text{Cov}(G_{i_{1}j_{1}},G_{i_{2}j_{2}})=\frac{1}{n}\sum_{k=1}^{n}\text{Cov}(Z_{ki_{1}}Z_{kj_{1}},Z_{ki_{2}}Z_{kj_{2}})$, $i_{1},i_{2},j_{1},j_{2}=1,\cdots,p$.
Propositions 1 and 2 follow as direct corollaries from Qiu, Li and Yao (2023) and Chernozhuokov et al. (2022), and we omit the proofs here. These propositions demonstrate that statistics based on the Frobenius and supremum norms lack the second-moment universality as the operator norm. Consequently, tests relying on the Frobenius and supremum norms require normalization through fourth-moment estimates of $\bm{X}$, a process more complex than estimating the covariance structure itself. This complexity arises because the Frobenius and supremum norms reduce the covariance matrices to vector forms, calculating the $L^{2}$ and $L^{\infty}$ norms, respectively. As a result, deriving the asymptotic distribution necessitates the second moment of the sample covariance, equivalent to the fourth moment of $\bm{X}$. Thus, the operator norm emerges as a more easy-to-implement option for covariance testing.
@@ 5.1 Simulation
In this subsection, we evaluate the empirical size and power of the proposed universal bootstrap for the operator norm-based statistics $T$ in (1) (Opn) and the combined statistics $T^{\text{Com}}$ in (9) (Com). These are compared with the operator norm-based statistic $T^{\text{Roy}}$ in (3) (Roy), the Frobenius norm-based linear spectral statistic $T^{\text{F},1}$ from Qiu, Li and Yao (2023) (Lfn), the debiased Frobenius norm-based U statistic $T^{\text{F},2}$ from Cai and Ma (2013) (Ufn), and the supremum norm-based statistic $T^{\text{sup}}$ inspired by Cai and Jiang (2011); Cai, Liu and Xia (2013) (Supn). The expressions for these statistics are provided in Section 12 of the Supplement.
To evaluate the robustness of our method, we generate $n$ independent, non-identically distributed samples of $p$-dimensional random vectors with zero means and covariance matrix $\bm{\Sigma}$. We set sample sizes $n=100$ and $n=300$ and vary the dimension $p$ over $100$, $200$, $500$, $1000$, and $2000$ to cover both proportional and ultra-high-dimensional cases. For each configuration, we calculate the average number of rejections based on $2000$ replications, with the universal bootstrap procedure using $1000$ resamplings. The significance level for all tests is fixed at $0.05$. We consider three different structures for the null covariance matrix $\bm{\Sigma}_{0}=(\sigma_{0,ij})_{p\times p}$:
@@
(b). Block diagonal. Let $K=\lfloor p/10\rfloor$ be the number of blocks, with each diagonal block being $0.55\bm{1}_{10}^{\ }\bm{1}_{10}^{T}+0.45\bm{I}_{10}$.
@@
These structures have been previously analyzed in Cai, Liu and Xia (2013); Zhu et al. (2017); Yu, Li and Xue (2024) for models (a) and (b), and in Cai, Liu and Xia (2013); Yu, Li and Xue (2024) for model (c), highlighting the broad applicability of our approach. To assess the empirical size, we examine the following distributions of $\bm{X}=\bm{\Sigma}^{\frac{1}{2}}\bm{Z}$:
(1). Gaussian distribution: $\{Z_{ij},\ i=1,\cdots,n;j=1,\cdots,p\}$ are i.i.d. standard normal random variables.
(2). Uniform and t-distributions: $\{Z_{ij},\ i=1,\cdots,\lfloor n/2\rfloor;j=1,\cdots,p\}$ are i.i.d. normalized uniform random variables on $[-1,1]$, while $\{Z_{ij},\ i=\lfloor n/2\rfloor+1,\cdots,n;j=1,\cdots,p\}$ are i.i.d. normalized t random variables with degree of freedom $12$.
(3). Gaussian and uniform distributions: $\{Z_{ij},\ i=1,\cdots,\lfloor n/2\rfloor;j=1,\cdots,p\}$ are i.i.d. standard normal random variables, while $\{Z_{ij},\ i=\lfloor n/2\rfloor+1,\cdots,n;j=1,\cdots,p\}$ are i.i.d. normalized uniform random variables on $[-1,1]$.
(4). Gaussian and t-distributions: $\{Z_{ij},\ i=1,\cdots,\lfloor n/2\rfloor;j=1,\cdots,p\}$ are i.i.d. standard normal random variables, while $\{Z_{ij},\ i=\lfloor n/2\rfloor+1,\cdots,n;j=1,\cdots,p\}$ are i.i.d. normalized t random variables with degree of freedom $12$.
Common covariance test cases often consider the Gaussian and uniform distributions Chen, Qiu and Zhang (2023), while the t-distribution with $12$ degrees of freedom is examined in Cai, Liu and Xia (2013); Zhu et al. (2017). All distributions are standardized to zero mean and unit variance. Except for the Gaussian baseline, the other cases share identical covariance matrices but differ in distribution, creating challenging yet crucial scenarios for testing covariance structures without imposing additional distributional assumptions. These cases reflect the broad applicability of our universality result.
Table 1: The empirical sizes of the supremum, the Frobenius, and the operator norm tests with a significance level $0.05$ for data with Gaussian distribution and the t and uniform distribution and the combinations of the sample size $n$, the dimension $p$, and the covariance matrix $\bm{\Sigma}_{0}$.

### v2

@@ 4 Application to covariance inference
$\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma}),\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)\ \|\bm{A}\|_{S_{1}}\bigg{]}\bigg{|}\bm{Y}^{1},\cdots,\bm{Y}^{B}\bigg{)}-(1-\alpha)\bigg{|}\lesssim\rho_{n}(\bm{\Sigma}_{0})+$ $d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))+B^{-\frac{1}{2}}.$
Here the Schatten $1-$norm of matrix $\bm{A}$ is defined as $\|\bm{A}\|_{S_{1}}=\sum_{i=1}^{p}\sigma_{i}(\bm{A})$. As a special case, we also have simultaneous confidence intervals for $\bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}$
$\sup_{0\leq\alpha\leq 1}\bigg{|}\mathbb{P}\bigg{(}\forall\bm{c}_{1},\bm{c}_{2}\in\mathbb{R}^{p},\ \bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}\in\bigg{[}\bm{c}_{1}^{T}\bm{\hat{\Sigma}}\bm{c}_{2}-\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma}),\widehat{\text{Sp}}(\bm{\Sigma})}$ $(\alpha)\ \|\bm{c}_{1}\|\cdot\|\bm{c}_{2}\|,\ \bm{c}_{1}^{T}\bm{\hat{\Sigma}}\bm{c}_{2}+$ (27)
@@
$\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma}),\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)\ \|\bm{c}_{1}\|\cdot\|\bm{c}_{2}\|\bigg{]}\bigg{|}\bm{Y}^{1},\cdots,\bm{Y}^{B}\bigg{)}-(1-\alpha)\bigg{|}$ $\lesssim\rho_{n}(\bm{\Sigma}_{0})+d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))+B^{-\frac{1}{2}}.$
Theorem 4.2 establishes sharp uniform simultaneous confidence intervals for both $\langle\bm{A},\bm{\Sigma}\rangle$ and $\bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}$. Specifically, intervals (26) and (27) are sharp, as their probability converges uniformly to $1-\alpha$, rather than no less than $1-\alpha$. Constructing confidence intervals for all linear combinations of $\bm{\Sigma}$ is considerably more challenging than for individual entries. For the mean vector case, achieving confidence intervals for each entry requires only $\ln p=o(n^{a})$ for some $a>0$, see Chernozhukov et al. (2013, 2017, 2023). In contrast, confidence intervals for all linear combinations of the mean vector require $p=o(n)$, see Zhai (2018); Xu et al. (2019); Fang & Koike (2024). Remarkably, we construct sharp simultaneous confidence intervals for all $\langle\bm{A},\bm{\Sigma}\rangle$ even as $p/n$ converges or diverges to infinity.
@@ 5.1 Simulation
In this subsection, we numerically analyze the performance of the universal bootstrap method in high-dimensional covariance tests. We evaluate the empirical size and power of the universal bootstrapped operator norm-based statistics $T$ in (1) (Opn) and the combined statistics $T^{\text{Com}}$ in (5) (Com). These are compared with the operator norm-based statistic $T^{\text{Roy}}$ in (3) (Roy), the Frobenius norm-based linear spectral statistic $T^{\text{F},1}$ from Qiu et al. (2023) (Lfn), the debiased Frobenius norm-based U statistic $T^{\text{F},2}$ from Cai & Ma (2013) (Ufn), and the supremum norm-based statistic $T^{\text{sup}}$ inspired by Cai & Jiang (2011); Cai et al. (2013) (Supn). The expressions for these statistics are provided in the Supplement.
To evaluate the robustness of our method, we generate $n$ independent, non-identically distributed samples of $p$-dimensional random vectors with zero means and covariance matrix $\bm{\Sigma}$. We set sample sizes $n=100$ and $n=300$ and vary the dimension $p$ over $100$, $200$, $500$, $1000$, and $2000$ to cover both proportional and ultra-high-dimensional cases. For each configuration, we calculate the average number of rejections based on $2000$ replications, with the universal bootstrap procedure using $1000$ resamplings. The significance level is fixed at $0.05$. We consider three different structures for the null covariance matrix $\bm{\Sigma}_{0}=(\sigma_{0,ij})_{p\times p}$:
@@
(b). Block diagonal. Let $K=\lfloor p/10\rfloor$ be the number of blocks, with each block being $1$ on the diagonal and $0.55$ on the off-diagonal.
@@
These structures have been previously analyzed in Cai et al. (2013); Zhu et al. (2017); Yu, Li & Xue (2024) for models (a) and (b), and in Cai et al. (2013); Yu, Li & Xue (2024) for model (c), highlighting the broad applicability of our approach. To assess the empirical size, we examine the following distributions: Gaussian distribution, Uniform and t-distributions, Gaussian and uniform distributions, and Gaussian and t-distributions. Here the Uniform and t-distributions sets $\{Z_{ij},\ i=1,\cdots,\lfloor n/2\rfloor;j=1,\cdots,p\}$ to be i.i.d. normalized uniform random variables on $[-1,1]$, and $\{Z_{ij},\ i=\lfloor n/2\rfloor+1,\cdots,n;j=1,\cdots,p\}$ to be i.i.d. normalized t random variables with degree of freedom $12$. The other distributions are defined similarly. All distributions of $Z_{ij}$ are standardized to zero mean and unit variance. Common covariance tests often consider the Gaussian and uniform distributions (Chen et al., 2023), while the t-distribution with $12$ degrees of freedom is examined in Cai et al. (2013); Zhu et al. (2017). Except for the Gaussian baseline, the other cases share identical covariance matrices but differ in distribution, creating challenging scenarios for testing covariance. These reflect the broad applicability of our universality result.
Table 1 reports the empirical sizes of the supremum, the Frobenius, and the operator norm tests at the significance level $0.05$, sample size $n=300$ across various dimension $p$, covariance $\bm{\Sigma}_{0}$ and data distributions of Gaussian and the uniform and t-distributions. For results on $n=100$ and for Gaussian-uniform and Gaussian-$t$ distributions with $n=100,300$, see Table 5-10 in the Supplement. The three operator norm tests are performed using the proposed universal bootstrap procedure, while the supremum and Frobenius norm tests rely on their respective asymptotic distribution formulas. Both Table 1 and Table 5-10 in the Supplement show that the universal bootstrap effectively controls the size of all operator norm-based statistics at the nominal significance level across all tested scenarios.
Table 1: The empirical sizes of different norm tests with a significance level $0.05$, the sample size $n=300$ for data with Gaussian distribution and the t and uniform distribution and the combinations of the dimension $p$, and the covariance matrix $\bm{\Sigma}_{0}$.